# Code Interviews: Design and Evaluation of a More Authentic Assessment for Introductory Programming Assignments

### v1

@@ Abstract.
Generative artificial intelligence poses new challenges around assessment and academic integrity, increasingly driving introductory programming educators to employ invigilated exams often conducted in-person on pencil-and-paper. But the structure of exams often fails to accommodate authentic programming experiences that involve planning, implementing, and debugging programs with computer interaction.
In this experience report, we describe code interviews: a more authentic assessment method for take-home programming assignments. Through action research, we experimented with varying the number and type of questions as well as whether interviews were conducted individually or with groups of students. To scale the program, we converted most of our weekly teaching assistant (TA) sections to conduct code interviews on 5 major weekly take-home programming assignments. By triangulating data from 5 sources, we identified 4 themes. Code interviews (1) pushed students to discuss their work, motivating more nuanced but sometimes repetitive insights; (2) enabled peer learning, reducing stress in some ways but increasing stress in other ways; (3) scaled with TA-led sections, replacing familiar practice with an unfamiliar assessment; (4) focused on student contributions, limiting opportunities for TAs to give guidance and feedback.
We conclude by discussing the different decisions about the design of code interviews with implications for student experience, academic integrity, and teaching workload.
academic integrity; authentic assessment; introductory programming; oral examinations;
@@ 1. Introduction
Generative artificial intelligence (AI) present new opportunities and new challenges for teaching, learning, and assessment in introductory programming courses (Lau and Guo, 2023; Prather et al., 2023; Becker et al., 2023; Savelka et al., 2023). Generative AI can help students receive immediate, personalized, and helpful feedback to support their learning. But the very same ability to assist students can also be used to circumvent learning. The availability, accessibility, and relatively widespread adoption of generative AI by students has challenged traditional ideas about the effectiveness of take-home assignments. Programming educators “need to prepare for a world in which there is an easy-to-use widely accessible technology that can be utilized by learners to collect passing scores, with no effort whatsoever, on what today counts as viable programming knowledge and skills assessments” (Savelka et al., 2023).
In particular, a recent study of an LLM-first introductory programming course by Vadaparty et al. reported that, anecdotally, “students’ performance on code writing (from scratch) questions was slightly lower than past offerings, but their performance on code tracing and code reading questions was roughly the same” (Vadaparty et al., 2024). Prather et al. build on this result with their study that identified how several participants “were often able to recognize a correct solution, but unable to get there themselves”, concluding that “Students may feel as though they are learning even as GenAI tools replace critical thinking and problem solving for them” (Prather et al., 2024). This has the potential to increase inequities as “students who are already poised to succeed can leverage GenAI to accelerate, while struggling students may be hindered by using GenAI, leaving them with an illusion of competence” (Prather et al., 2024). Educators need reliable methods to evaluate student learning outcomes.
@@
While invigilated in-person pencil-and-paper exams remain a relatively reliable way to evaluate student learning outcomes, introductory programming courses often include different types of take-home work that resist strictly time-limited and computer-free evaluation. For example, educators may want to assess students’ abilities to not only solve many small problems, but also apply their skills to assemble larger programs. This format might be harder to squeeze into small, fixed exam periods without computer access for students to run and debug code. Additionally, just as there are equity implications for the use of generative AI in computing education (Prather et al., 2024), there are also equity implications for assessment policies as different decisions can impact student motivation, sense of belonging, and self-efficacy (Sherif et al., 2024). If authentic programming experiences are important learning objectives, how might we design a more robust assessment for take-home work while minimizing academic misconduct? How might we design more authentic and meaningful assessments that promote this goal of academic integrity while maintaining or increasing student motivation, sense of belonging, and self-efficacy? How might we achieve all of this without reinventing our existing courses and assignments?
@@
In this experience report, we describe code interviews: a method to evaluate student proficiency on programming-related learning outcomes using live oral assessments that extend students’ take-home programming assignments. We integrated our approach over two academic terms (quarters) in a large-enrollment undergraduate intermediate data programming course. Drawing on action research methodology, we improved our approach over the two quarters using data triangulated from 5 different sources: observations, semi-structured focus groups, surveys, small-group instructional diagnosis, and teaching assistant (TA) feedback in order to answer three research questions.
@@ 3. RQ3
Oral exams have long been used in education settings to assess student proficiency and learning outcomes. In recent years, there have been many reasons for adopting oral exams in science, engineering, and computing courses, such as:
  • increasing confidence in assessments particularly due to disruptions of global health or new technologies (Gardner and Giordano, 2023; Sabin et al., 2021; Lubarda et al., 2021),
  • improving emotional reactions and perceptions about assessment effectiveness (Sabin et al., 2021; Ohmann, 2019),
  • improving student motivation, stress level, overall course performance, and sense of belonging (Reckinger and Reckinger, 2022)
Results from these studies have been mixed. Depending on the particular approach taken, study population, and research methods, studies have varied in their ability to produce desirable educational outcomes. In undergraduate chemistry education, Gardner and Giordano report that “Students had an overwhelmingly positive response to the oral exam experience and recommended their continued use in spite of […] the stress and anxiety of verbal presentation and the depth of understanding required to answer questions verbally” (Gardner and Giordano, 2023). In undergraduate computing education, Sabin et al. found that student experiences varied by self-identified demographic characteristics, with women, Black or African American, Hispanic or Latino, and some subgroups of first-generation college students reporting higher stress levels before oral exams than their majority group peers (Sabin et al., 2021). The authors also identified similarly concerning patterns of inequity for overall course performance and sense of belonging. Despite these challenges, many authors still consider oral exams a promising assessment (Gardner and Giordano, 2023; Sabin et al., 2021; Lubarda et al., 2021; Theobold, 2021; Ohmann, 2019) with the potential to make assessment more authentic (Theobold, 2021) and promote academic integrity (Lubarda et al., 2021).
@@ 2. Code Interviews
Code interviews are a type of live oral assessment for programming. In the context of our course—an undergraduate intermediate data programming course enrolling about 300 students during winter quarter and 250 students during spring quarter—code interviews were offered only in-person and designed to evaluate students’ ability to explain and extend their weekly take-home programming work. As this required discussing their take-home programs, discussions were conducted around the due date for each of the 5 weekly programming assignments. These code interviews were conducted during TA-led sections, which we over-staffed to a ratio of 3 TAs per 30 student section. This allowed us to conduct 3 code interviews in parallel during each section, giving us space to experiment with variables such as the number of students who conducted a code interviews at once (group size) and the questions TAs asked during code interview (and, therefore, the expected duration of each interview).
@@ 2.1. Winter Quarter
$2.$ Spring Quarter
In the following spring quarter, the course staff sought to increase standardization by asking more specific questions. During spring quarter, code interviews focused on specific questions drawn from a selection of question types such as tracing a test case, comparing approaches, identifying an alternative approach, finding a bug, and writing code to solve an isomorphic problem. For each question, TAs were empowered to give each student one nudge question that would help them make progress. Students were given the question types each week in advance and evaluated on their ability to satisfactorily answer the questions. We experimented with group size between each week.
@@ 2.2.1. First Code Interview
The first code interview during spring quarter extended the weekly take-home programming assignment that reviewed basic Python functionality and programming concepts. This was conducted in groups of 2 students to better assess individual students while maintaining the social and comparative discussion aspect from winter quarter code interviews. During this code interview, we asked three types of questions to each student: tracing a test case, comparing approaches, and identifying an alternative approach.
@@
Identifying an alternative approaches required the student to figure out how to solve one of the problems from their take-home programming assignment in a different way. Scenarios included constraints such as solving the problem without any data structures, with a different data structure, etc. This type of question pushed students to think about different the advantages and disadvantages of different approach.
These three types of questions represented a good balance of different skills that we wanted to assess, but we faced several issues. There were too many questions for the time frame we prepared, so TAs and students felt rushed. Student work also often converged on the same approaches, so comparison between approaches was limited. Finally, students found it difficult to develop a new approach under the constraints of the live code interview. The next code interview was adjusted according to these experiences.
@@ 2.2.2. Second Code Interview
The second code interview extended the corresponding assignment on data manipulation and analysis. To alleviate time pressure, we moved from three questions to two questions with the same time limit. The two types of questions were tracing a test case (as before, intended to help warm up) and finding a bug.
For finding a bug, given a buggy solution to a part of the weekly assignment (and, sometimes, the result of running it), students were tasked to figure out the bug. Depending on the particular question, the output was sometimes given as a nudge and sometimes given originally with the problem itself. This question evaluated students’ debugging skills as well as their understanding of their own code. When the output was not given, this question tested students’ understanding of the problem and code comprehension in order to to identify the approach and the potential bug within it.
@@ 2.2.3. Third Code Interview
The third code interview extended the corresponding assignment on data visualization. This code interview was conducted entirely individually. TAs asked two types of questions: finding a bug and writing code to solve an isomorphic problem. Isomorphic problems are very similar to problems in the weekly take-home programming assignment, designed so that students could change just a few small parts of their assignment solution to solve the isomorphic problem. TAs evaluated students on not only their solution to the isomorphic problem but also their explanation of their work.
The main difficulty with this structure was that the isomorphic question took a long time since students had to read the entire question again and then develop a solution, including dealing with any errors along the way. We decided that the next time we asked an isomorphic question, it would be the only question we asked.
@@ 2.2.4. Fourth Code Interview
The fourth code interview extended the corresponding assignment on software engineering for data science. Since the engineering task involved implementing a complicated algorithm, TAs asked two types of questions: tracing their solution and finding a bug. We decided to conduct this code interview in pairs because we were also getting feedback from students that the interviews were pretty stressful. We began liking the ability to ask different types of questions each week: as assignments changed from week to week, code interviews also benefited from changes between each week.
@@ 2.2.5. Fifth Code Interview
The fifth and final code interview extended the corresponding assignment on geospatial data visualization. For this interview, we wanted to focus on writing code to solve an isomorphic problem, which (as decided earlier) would be the only type of question we asked. Drawing from the third code interview where we last asked students to write code to solve an isomorphic problem, we wanted to run this fifth and final code interview the same way. However, taking into account feedback from students that they appreciated having a partner, we ultimately conducted this code interview in pairs but with separate questions and limited collaboration. We hypothesized that students might be able to gain some comfort from working alongside another student even if collaboration opportunities were limited.
@@ 3. Methods
We evaluated code interviews using weekly student surveys, observations of code interviews, semi-structured student focus groups, a class-wide small-group instructional diagnosis, as well as a few individual TA interviews and group TA discussions. We used inductive thematic analysis to identify key themes and patterns in the data.
Students completed a pre-survey, a post-survey, and 4 weekly surveys in between each of the 5 weekly take-home programming assignments. The pre-survey included questions about students’ comfort and familiarity with code interviews, as well as their sense of its importance. The post-survey included questions about students’ perception of code interviews: what skills they believed they gained, self-confidence in their code interview skills, and how code interviews have change their perception on computing. At the end of the post-survey, we also asked students about demographic information to determine if there were disparities in experiences between students with different social identities; however, in this study, we did not utilize this information. For the weekly survey, we asked students the same set of three questions each week:
@@ 4. Results
Code interviews enabled students to learn from each other. One student said, “When I ran [into] an issue—or some problem someone else ran into—they [peers] explained how they solved it” (spring focus group). Because students are given time to think about how to answer questions, code interviews also provided an opportunity for them to discuss the assignment and problems that occurred when completing it, uncovering common mistakes and fixes. During the spring small-group instructional diagnosis, students mentioned that seeing how their peers solved problems helped them understand the concepts. Another student noted that “it’s effective if you have two people sharing their inputs and outputs. As long as it’s constructive criticism, there [are] always ideas that you have never thought of unless someone else says it, so it helps to have another input” (spring focus group). Our section observations corroborated this perspective. In addition to discussing their approaches during the code interview, students also had a chance to discuss their work with each other during the remainder of the scheduled section time while TAs were busy helping other groups conduct their code interviews.
Particularly in an assessment activity that might otherwise heighten student stress, conducting code interviews with groups of students may reduce stress compared to individual, one-on-one code interviews with a TA. Some students particularly appreciated group code interviews because they could discuss their solutions with peers to improve the quality of their weekly programming assignment submission. Although the benefits of peer learning may have occurred most in group code interviews, even individual code interviews may have still enabled peer learning. During spring quarter sections, we observed students talking among themselves to help improve each other’s solutions and prepare for code interviews even during weeks when code interviews were conducted individually. Alternating between group code interviews, individual code interviews, and collaborative section practice activities may provide a diversity of ways for students to teach and learn from each other.
@@
$4.$ Code interviews focused on student contributions, limiting opportunities for TAs to give guidance and feedback

### v2

@@ Abstract.
Generative artificial intelligence poses new challenges around assessment, increasingly driving introductory programming educators to employ invigilated exams. But exams do not afford more authentic programming experiences that involve planning, implementing, and debugging programs with computer interaction. In this experience report, we describe code interviews: a more authentic assessment method for take-home programming assignments. Through action research, we experimented with varying the number and type of questions as well as whether interviews were conducted individually or with groups of students. To scale the program, we converted most of our weekly teaching assistant (TA) sections to conduct code interviews on 5 major weekly take-home programming assignments. By triangulating data from 5 sources, we identified 4 themes. Code interviews (1) pushed students to discuss their work, motivating more nuanced but sometimes repetitive insights; (2) enabled peer learning, reducing stress in some ways but increasing stress in other ways; (3) scaled with TA-led sections, replacing familiar practice with an unfamiliar assessment; (4) focused on student contributions, limiting opportunities for TAs to give guidance and feedback. We conclude by discussing the different decisions about the design of code interviews with implications for student experience, academic integrity, and teaching workload.
academic integrity; authentic assessment; code interviews; introductory programming; large classes; teaching assistants; oral exams
@@ 1. Introduction
Generative artificial intelligence (AI) present new opportunities and new challenges for teaching, learning, and assessment in introductory programming courses (Lau and Guo, 2023; Prather et al., 2023; Becker et al., 2023; Savelka et al., 2023; Novak and Ohmann, 2023). The same capability that enables students to receive immediate, personalized, and helpful feedback to support their learning can also be used to circumvent learning. The availability and accessibility of generative AI poses a challenge to the accuracy of take-home assessment (Novak and Ohmann, 2023). Educators “need to prepare for a world in which there is an easy-to-use widely accessible technology that can be utilized by learners to collect passing scores, with no effort whatsoever, on what today counts as viable programming knowledge and skills assessments” (Savelka et al., 2023).
@@
While invigilated in-person pencil-and-paper exams remain a relatively reliable way to evaluate student learning outcomes, programming courses often include take-home work that resist time-limited and computer-free assessment. For example, educators may want to assess students’ abilities to not only solve many small problems, but also assemble larger programs. This format might be harder to squeeze into small, fixed exam periods without computer access for students to run and debug code. Additionally, just as there are equity implications for the use of generative AI in computing education (Prather et al., 2024), there are also equity implications for assessment policies as different decisions can impact student motivation, sense of belonging, and self-efficacy (Sherif et al., 2024). If authentic programming experiences are important learning objectives, how might we design a more robust assessment for take-home work while minimizing academic misconduct? How might we design more authentic and meaningful assessments that promote this goal of academic integrity while maintaining or increasing student motivation, sense of belonging, and self-efficacy? How might we achieve all of this without reinventing our existing courses and assignments?
@@
In this experience report, we describe code interviews: a method to evaluate student proficiency on programming learning outcomes using live oral assessments about students’ take-home work. We integrated our approach over two academic terms (quarters) in a large-enrollment undergraduate intermediate data programming course (an introductory python course taught in the context of data science). Drawing on action research methodology, we improved our approach over the two quarters using data triangulated from 5 different sources: observations, semi-structured focus groups, surveys, small-group instructional diagnosis, and teaching assistant (TA) feedback in order to answer three research questions.
@@ 3. RQ3
Oral exams have long been used in education settings to assess student proficiency and learning outcomes (Gharibyan, 2005). In recent years, there have been many reasons for adopting oral exams in science, engineering, and computing courses, such as:
  • increasing confidence in assessments particularly due to disruptions of global health or new technologies (Gardner and Giordano, 2023; Sabin et al., 2021; Lubarda et al., 2021);
  • improving perceptions about assessment effectiveness (Sabin et al., 2021; Ohmann, 2019; Gharibyan, 2005) as well as motivation, stress level, overall course performance, and sense of belonging (Reckinger and Reckinger, 2022);
  • emphasizing communication skills and assess student learning outcomes in light of generative AI (Novak and Ohmann, 2023).
Results from these studies have been mixed. Depending on the approach taken, study population, and research methods, studies have varied in their ability to produce desirable educational outcomes. In undergraduate chemistry education, Gardner and Giordano report that “Students had an overwhelmingly positive response to the oral exam experience and recommended their continued use in spite of […] the stress and anxiety of verbal presentation and the depth of understanding required to answer questions verbally” (Gardner and Giordano, 2023). In undergraduate computing education, Sabin et al. found that student experiences varied by self-identified demographic characteristics, with women, Black or African American, Hispanic or Latino, and some subgroups of first-generation college students reporting higher stress levels before oral exams than their majority group peers (Sabin et al., 2021). The authors also found similarly concerning patterns of inequity for overall course performance and sense of belonging. Novak and Ohmann reiterate two problems about oral exams: “(1) planning the timing and logistics and (2) evaluating students responses in an ethical and unbiased way” (Novak and Ohmann, 2023). Despite these challenges, many authors still consider oral exams a promising assessment (Gardner and Giordano, 2023; Sabin et al., 2021; Lubarda et al., 2021; Theobold, 2021; Ohmann, 2019) with the potential to make assessment more authentic (Theobold, 2021) and promote academic integrity (Lubarda et al., 2021).
@@ 2. Code Interviews
Code interviews are a type of live oral assessment for programming. In the context of our course—an intermediate Python data science course at a large research-intensive university typically enrolling over 200 students—code interviews were offered only in-person and designed to evaluate students’ ability to explain and extend their weekly take-home programming work. As this required discussing their take-home work, discussions were conducted on the due date for each of the 5 weekly programming assignments during the weekly 50-minute TA-led section. In order to scale code interviews, we over-staffed sections to a ratio of 3 TAs per 30 student section in order to host 3 five-minute code interviews in parallel during each section. This over-staffed section structure gave us space to experiment with research variables such as the number of students who conducted a code interviews at once (group size) and the type of questions TAs asked during code interview (and, therefore, the expected duration of each interview).
@@ 2.1. Winter Quarter
Code interviews were evaluated by TAs using a discrete, 4-level rubric from greatest proficiency to least proficiency: Exemplary, Satisfactory, Approaching, and Unassessable.
Across the entire winter quarter, 76% of all code interviews conducted received Exemplary marks while the remaining 24% received Satisfactory marks. Assignment completion rates as well as code interview participation rates were very high as they were both required by the course.
2.2. Spring Quarter
In the following spring quarter, the course staff wanted to increase the “signal-to-noise ratio” of code interviews by asking more specific questions and assessing students in smaller groups (or even individually). Spring quarter code interviews focused on specific questions drawn from a selection of question types: tracing a test case, comparing approaches, identifying an alternative approach, finding a bug, and writing code to solve an isomorphic problem. Each week, the course staff announced the types of questions that would be asked. During the code interview, TAs would choose questions from a question pool and evaluate the student’s live oral response to it. If, after a short while, the student seemed stuck, TAs were empowered to give each student one nudge question to help them make progress. Proficiency was determined considering both the correctness of the answer and effectiveness of the explanation.
Across the entire spring quarter, 30% of all code interviews received Exemplary marks, 63% received Satisfactory marks, and the remaining 7% received Approaching marks. Assignment completion rates as well as code interview participation rates were very high as they were both required by the course.
@@ 2.2.1. First Code Interview
The first code interview during spring quarter extended the weekly take-home programming assignment that reviewed Python programming concepts. This was conducted in groups of 2 students to better assess individual students while maintaining the social and comparative discussion aspect from winter quarter code interviews. During this code interview, we asked three types of questions to each student: tracing a test case, comparing approaches, and identifying an alternative approach.
@@
Identifying an alternative approach required the student to solve one of the problems from their take-home programming assignment in a different way. Scenarios included constraints such as solving the problem without any data structures, with a different data structure, etc. This type of question pushed students to think about different the advantages and disadvantages of different approach.
These question types provided excellent coverage of the learning outcomes, but we faced several difficulties. There were too many questions for the time frame we prepared, so TAs and students felt rushed. Student work also often converged on the same approaches, so comparison between approaches was limited. Finally, students found it difficult to develop a new approach in front of a TA. We adjusted the next code interview according to these experiences.
@@ 2.2.2. Second Code Interview
The second code interview extended the corresponding assignment on statistical data analysis. To alleviate time pressure, we moved from three questions to two questions with the same time limit. The two types of questions were tracing a test case (as before, intended to help warm up) and finding a bug.
Finding a bug involved identifying the defects in a TA-provided sample solution to the weekly programming assignment. Depending on the particular question, the output of running the defective code was sometimes given as a nudge and sometimes given initially with the problem itself. This question evaluated students’ debugging skills as well as their understanding of their own code. When the output was not given, this question tested students’ understanding of the problem and code comprehension in order to to identify the approach and the potential bug within it.
@@ 2.2.3. Third Code Interview
The third code interview extended the corresponding assignment on data visualization. This code interview was conducted entirely individually: TAs met with students one-on-one. TAs asked two types of questions: finding a bug and writing code to solve an isomorphic problem.
Isomorphic problems are very similar to problems in the weekly take-home programming assignment—designed so that students only needed to modify a few small parts of their programming assignment in order to solve the new isomorphic problem. TAs evaluated students on not only their solution to the isomorphic problem but also their explanation of their work.
The main difficulty with this structure was that the isomorphic question took a long time since students had to read the entire question again and then develop a solution, including dealing with any errors along the way. Time again became an issue, particularly since these code interviews were conducted individually. We decided that the next time we asked an isomorphic question to students individually, it would be the only question we asked.
@@ 2.2.4. Fourth Code Interview
The fourth code interview extended the corresponding assignment on software engineering. Since the engineering task involved implementing a complicated algorithm, TAs asked two types of questions: tracing their solution and finding a bug. We decided to conduct this code interview in pairs in response to feedback from students that interviews were stressful. We appreciated asking different types of questions each week: as assignments changed from week to week, code interview questions also changed to reflect them.
@@ 2.2.5. Fifth Code Interview
The fifth and final code interview extended the corresponding assignment on geospatial data visualization. For this interview, we wanted to focus on writing code to solve an isomorphic problem individually, which (as decided earlier) would be the only type of question we asked. Drawing from the third code interview where we last asked students to write code to solve an isomorphic problem, we wanted to run this fifth and final code interview the same way. However, taking into account feedback from students that they appreciated having a partner, we ultimately conducted this code interview in pairs but with separate questions and limited collaboration. We hypothesized that students might be able to gain some comfort from working alongside another student even if collaboration opportunities were limited.
@@ 3. Methods
We evaluated code interviews using weekly student surveys, observations of code interviews, semi-structured student focus groups, a class-wide small-group instructional diagnosis, as well as a few individual TA interviews and group TA discussions. We identified key themes in the data using inductive thematic analysis. None of the research instruments listed above were required, so student participation in the research evaluation was entirely optional.
Students were asked to complete a pre-survey, a post-survey, and 4 weekly surveys in between each of the 5 take-home programming assignments. The pre-survey included questions about students’ comfort and familiarity with code interviews, as well as their sense of its importance. The post-survey included questions about students’ perception of code interviews: what skills they believed they gained, self-confidence in their code interview skills, and how code interviews have change their perception on computing. At the end of the post-survey, we also asked students about demographic information to determine if there were disparities in experiences between students with different social identities; however, in this study, we did not utilize this information. For the weekly survey, we asked students the same set of three questions each week:
@@
3.1. Student Demographics
Across both quarter, the self-reported student demographics were similar. Approximately 65% of students identified as Asian, 25% as White, and around 4% as each of Black or Hispanic/Latino. Students of all other races collectively accounted for less than 3%. For gender representation, there was a slightly higher proportion of women compared to men, less than 1% identified as non-binary, and a small remainder preferred not to say. About 40% of the class were English Language Learners, and roughly 10% reported having disabilities.
Introductory and intermediate programming courses offered at our university have large enrollments typically between 200 and 800 students. For early-career undergraduate students, code interviews may have been the most personalized interactions that they have had with computing course staff about their programming work.
@@ 4. Results
Code interviews enabled students to learn from each other. One student said, “When I ran [into] an issue—or some problem someone else ran into—they [peers] explained how they solved it” (spring focus group). Because students are given time to think about how to answer questions, code interviews also provided an opportunity for them to discuss the assignment and problems that occurred when completing it, uncovering common mistakes and fixes. During the spring small-group instructional diagnosis, students mentioned that seeing how their peers solved problems helped them understand the concepts. Another student noted that “it’s effective if you have two people sharing their inputs and outputs. As long as it’s constructive criticism, there [are] always ideas that you have never thought of unless someone else says it, so it helps to have another input” (spring focus group). Our section observations corroborated this perspective. In addition to discussing their approaches during the code interview, students also had a chance to discuss their work with each other during the remainder of the scheduled section time while TAs were busy interviewing other groups.
Conducting code interviews with groups of students may reduce stress compared to individual, one-on-one code interviews with a TA. Some students particularly appreciated group code interviews because they could discuss their solutions with peers to improve the quality of their weekly programming assignment submission. Although the benefits of peer learning may have occurred most in group code interviews, even individual code interviews may have still enabled peer learning. During spring quarter sections, we observed students talking among themselves to help improve each other’s solutions and prepare for code interviews even during weeks when code interviews were conducted individually. Alternating between group code interviews, individual code interviews, and collaborative section practice activities may provide a diversity of ways for students to teach and learn from each other.
@@
4.4. Code interviews focused on student contributions, limiting opportunities for TAs to give guidance and feedback